OpenStudy (anonymous):

you toss 3 coins. what is the probability that you get exactly two heads , given that you at least get one head?

OpenStudy (anonymous):

wouldnt it be 2/3?

OpenStudy (anonymous):

yes it is..i got it after. thank you

OpenStudy (anonymous):

your welcome!

OpenStudy (zarkon):

why 2/3?

OpenStudy (anonymous):

because there is 3 coins and only 2 possibilities?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

the answer should be 3/7

OpenStudy (anonymous):

ohhh

OpenStudy (zarkon):

P(2H|at least one head) =P(2H,at least one head)/P(at least one head) =P(2H)/[1-P(3 tails)]=(3/8)/(1-1/8)=(3/8)/(7/8)=3/7

OpenStudy (anonymous):

too confusing for me

OpenStudy (zarkon):

use the definition of conditional probability \[P(A|B)=\frac{P(A,B)}{P(B)}\]

OpenStudy (zarkon):

list out all the possibilities HHH HHT HTH THH HTT THT TTH TTT only these have at least one head (there are 7 of them) HHH HHT HTH THH HTT THT TTH of these only 3 of them have 2 heads

OpenStudy (anonymous):

ok....too confusing for me, but i kinda get it a little.